Which point is on the graph of f(x) = –3sin x + 14?

A. (1.5π, 8)
B. (1.5π,12)
C. (π, 14)
D. (π, 17)
1
Expert's answer
2013-08-05T08:26:52-0400
To find out if the point is on the graph we should substitute its coordinates into the given equation:
A: 8=-3sin (1.5π)+14=-3*(-1)+14=17 false
B: 12=-3sin (1.5π)+14=-3*(-1)+14=17 false

C: 14=-3sin (π)+14=14 true
D: 17=-3sin (π)+14=14 false
Thus, the rigth answer is C.